Shift Summary for day shift, 11/04/99

Jianbo Wang (jwang@physics.arizona.edu)
Thu, 4 Nov 1999 14:34:18 -0700 (MST)

Shift Summary for day shift, 11/04/99

On Shift: Leo, Jianbo

Hours of beam available at all intensities: 0.0
Hours of beam available at nominal intensity: 0.0
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

tevatron still down (cyro problems).
========================================================================
Runs ended this shift:

Tapes brought to FCC this shift:
(including tapes for online split)

none
========================================================================
Accesses during shift:

none

========================================================================
Overview of shift:

JasonL tried to fix the KQ problem which caused the ktevana error of
unknown mod type

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

7:25 Shift crew is shocked to discovered that we have
run completely out of fireballs!

7:36 A lot of ktevana errors were caused by failed
readout (Not the KQ, but the blossom) for crate 4(3Y)
Jason worked on it, but he was too tired to finish
it. So we stopped the trigger to avoid audio
alarm of ktevana error.

13:50 Gas check done.

JasonH worked on Crate 53 problem.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Beam is supposed to come at 4:40pm.